COLLISION-INDUCED INFRARED ABSORPTION BY COLLISIONAL COMPLEXES IN DENSE HYDROGEN-HELIUM GAS MIXTURES AT THOUSANDS OF KELVIN

\underline{MARTIN ABEL}, LOTHAR FROMMHOLD, Department of Physics, The University of Texas at Austin, Austin, TX 78712; XIAOPING LI, KATHARINE L. C. HUNT, Department of Chemistry, Michigan State University, East Lansing, MI 48824.

The interaction-induced absorption by collisional pairs of H_{2} molecules is an important opacity source in the atmospheres of the outer planets and cool stars ^{a}. The emission spectra of cool white dwarf stars differ significantly in the infrared from the expected blackbody spectra of their cores, which is largely due to absorption by collisional H_{2}–H_{2}, H_{2}–He, and H_{2}–H complexes in the stellar atmospheres. Using quantum-chemical methods we compute the atmospheric absorption from hundreds to thousands of kelvin ^{b}. Laboratory measurements of interaction-induced absorption spectra by H_{2} pairs exist only at room temperature and below. We show that our results reproduce these measurements closely ^{c}, so that our computational data permit reliable modeling of stellar atmosphere opacities even for the higher temperatures ^{d}.

aL. Frommhold, Collision-Induced Absorption in Gases, Cambridge University Press, Cambridge, New York, 1993 and 2006

bXiaoping Li, Katharine L. C. Hunt, Fei Wang, Martin Abel, and Lothar Frommhold, Collision-Induced Infrared Absorption by Molecular Hydrogen Pairs at Thousands of Kelvin, International Journal of Spectroscopy, vol. 2010, Article ID 371201, 11 pages, 2010. doi: 10.1155/2010/371201

cM. Abel, L. Frommhold, X. Li, and K. L. C. Hunt, Collision-induced absorption by H_{2} pairs: From hundreds to thousands of Kelvin, J. Phys. Chem. A, published online, DOI: 10.1021/jp109441f

dL. Frommhold, M. Abel, F. Wang, M. Gustafsson, X. Li, and K. L. C. Hunt, ”Infrared atmospheric emission and absorption by simple molecular complexes, from first principles”, Mol. Phys. 108, 2265, 2010